DISCOVERY OF THE ELECTRONIC SPECTRA OF HPS AND DPS

\underline{ROBERT A. GRIMMINGER}, JIE WEI, BLAINE ELLIS and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY, 40506; ZHONG WANG and TREVOR SEARS, Department of Chemistry, Brookhaven National Lab, Upton, NY 11973.

The hitherto unknown electronic spectrum of the closed shell transient molecule HPS has been observed in the 685 - 846 nm region by laser-induced fluorescence and single vibronic level emission techniques. HPS (and DPS) were produced in a pulsed electric discharge jet using a precursor mixture of 3% PH_{3} and 1% H_{2}S (or PD_{3} and D_{2}S) in high pressure argon. The weak set of observed bands are assigned to the Ã^{1}A′′ − X̃^{1}A′ electronic transition on the basis of chemical evidence, isotope shifts and the correspondence of the vibrational frequencies, excitation energy, and band contours with predictions based on our own high level ab initio calculations. Theory predicts that the HPS bond angle decreases on electronic excitation, contrary to expectations based on Walsh diagrams.